A metra commuter train leaves Union Station in Chicago at 12 noon. Two hours later, an amtrak train leaves on the same track, traveling at an average speed that is 50 mph faster that the metra train. At 3 pm, the amtrak train is 10 miles behind the commuter train. How fast is each going?
And how did i conclude this answer?

Answer by checkley77(12569)

D=RT
FOR THE METRA D=R*3
FOR THE AMTRAX D-10=(R+50)*1
R3=R+50+10
3R-R=60
2R=60
R=60/2
R=30 MPH FOR THE MATRA TRAIN.
30+50=80 MPH FOR THE AMTRAX TRAIN.
PROOF:
3*30=30+50+10
90=90